# ADSeeker: A Knowledge-Infused Framework for Anomaly Detection and Reasoning

### v1

@@ Abstract
Automatic vision inspection holds significant importance in industry inspection. While multimodal large language models (MLLMs) exhibit strong language understanding capabilities and hold promise for this task, their performance remains significantly inferior to that of human experts. In this context, we identify two key challenges: (i) insufficient integration of anomaly detection (AD) knowledge during pre-training, and (ii) the lack of technically precise and context-aware language generation for anomaly reasoning. To address these issues, we propose ADSeeker, an anomaly task assistant designed to enhance inspection performance through knowledge-grounded reasoning. ADSeeker leverages a curated visual document knowledge base, SEEK-MVTec&VisA (SEEK-M&V), which we construct to address the limitations of existing resources that rely solely on unstructured text. SEEK-M&V includes semantic-rich descriptions and image-document pairs, enabling more comprehensive anomaly understanding. To effectively retrieve and utilize this knowledge, we introduce the Query Image-Knowledge Retrieval-Augmented Generation (Q2K RAG) framework. To further enhance the performance in zero-shot anomaly detection (ZSAD), ADSeeker leverages the Hierarchical Sparse Prompt mechanism and type-level features to efficiently extract anomaly patterns. Furthermore, to tackle the challenge of limited in industry anomaly detection (IAD) data, we introduce the largest-scale AD dataset, Multi-type Anomaly (MulA), encompassing 72 multi-scale defect types across 26 Categories. Extensive experiments show that our plug-and-play framework, ADSeeker, achieves state-of-the-art zero-shot performance on several benchmark datasets.
@@ 1 Introduction
Figure 1: Industrial and medical image anomaly reasoning results of ADSeeker.
@@
Multimodal Large Language Models (MLLMs), such as GPT-4o [et al, 2024] and LLaVA [Li et al., 2024], have shown significant promise across various intellectual human tasks, including writing, programming, and Industry Anomaly Detection (IAD). In IAD tasks, anomaly reasoning is distinct from anomaly detection. Anomaly reasoning necessitates a finer level of granularity and logical reasoning in the model’s responses. This distinction is vital in identifying the roles of a quality inspector versus a domain expert. Recent studies have explored the use of MLLMs for IAD tasks, including detection, description, analysis, etc., by training models such as AnomalyGPT [Gu et al., 2023] and Anomaly-OV [Xu et al., 2025], to incorporate domain-specific knowledge for improved anomaly reasoning. Although some models excel in zero-shot anomaly detection (ZSAD), part of them struggle to provide accurate defect analysis and detailed object descriptions during the testing phase when engaging in anomaly reasoning. It should be noted that training methods for IAD MLLMs may seriously compromise the pre-trained knowledge and generalization of the model and the lack of IAD training data can lead to overfitting in the content of the answers. To tackle these issues, we propose ADSeeker, which addresses the limitations of domain knowledge through the Query Image—Knowledge Retrieval-Augmented-Generation (Q2K RAG).
@@
ADSeeker achieves multimodal hybrid retrieval, seeking the knowledge document most relevant to the query image. Specifically, the query image and knowledge document are encoded into the Key Embedding and the Lock Embedding respectively. Subsequent multimodal retrieval is performed in the joint feature space between these representations. To compensate for the lack of knowledge base in IAD domain, we establish the first-ever knowledge base SEEK-M&V composed of visually rich industry anomaly documents. In SEEK-M&V, each knowledge document has specific reference pages containing the defect type and the defect analysis. Moreover, we utilize DeepSeek-R1 [et al, 2025] to generate semantic-rich descriptions and application scenario introductions to expand the content of SEEK-M&V. As the saying goes, “a picture is worth a thousand words.” SEEK-M&V enriches knowledge representation by preserving graphic information for each anomaly type, offering a valuable reference for anomaly reasoning. Compared to existing training-based approaches, it significantly improves both efficiency and effectiveness.
@@
Existing approaches have introduced learnable prompts to adapt models for ZSAD. However, object-level prompts (e.g., identifying objects such as “wood" or “leather") are incapable of providing detailed anomaly patterns, and publicly available AD datasets struggle with semantically aligning these patterns. To address these gaps, we introduce Multi-type Anomaly (MulA), the largest-scale dataset for ZSAD, which includes 72 defect types across 8 scenarios. The key features of MulA, encoded through the Q2K RAG framework, are clustered using centroid-based techniques aligned with predefined cluster centers. Additionally, ADSeeker utilizes a Hierarchical Sparse Prompt (HSP) mechanism to extract type-level features (e.g., distinguishing between “scratch" or “hole" anomalies) and reduce hallucinations caused by suspected anomaly regions.
@@
Extensive experiments validate the effectiveness and efficiency of ADSeeker. For example, anomaly reasoning experiments on MMAD [Jiang et al., 2024] benchmark demonstrate that ADSeeker outperforms other general models. We also compare ADSeeker with best-performing training-based methods in IAD domain. Additionally, experiments on 12 datasets from both industrial and medical sectors highlight ADSeeker’s superiority in ZSAD tasks. ADSeeker effectively extracts anomaly patterns at the type level, offering a novel approach to anomaly reasoning. To validate its effectiveness, we performed extensive tests to visualize the impact of type-level features across a range of products. We also conducted ablation studies and evaluated the performance-resource trade-offs within our ADSeeker.
@@ 2 Related Work
MLLM for industry anomaly detection (IAD). Recent advances have demonstrated the potential of MLLMs in IAD. For example, AnomalyGPT is directly fine-tuned on IAD datasets to inject domain knowledge into the base model. Moreover, to effectively extract detailed feature from abnormal samples, Myriad [Li et al., 2023] developed an Expert Perception module, which is designed to harness the prior knowledge of vision experts. However, these models are constrained by the scope of their training knowledge and exhibit limited domain generalization. Existing IAD approaches rarely become a trustworthy automatic vision inspection assistant [Lin et al., 2023, wen Dong et al., 2024, Yao et al., 2024, Chen et al., 2023, Liu et al., 2023a, Liu et al., 2024]. In contrast, building on the excellence of our ADSeeker framework, the anomaly reasoning performance of base models have been notably enhanced, as shown in Table 5.
Zero-Shot Anomaly Detection (AD). Recently, WinCLIP [Jeong et al., 2023] makes an early attempt at introducing VLMs into ZSAD tasks, which leverage CLIP [Radford et al., 2021] to compute the similarities between patch embeddings and textual embeddings. On this basis, AnomalyCLIP [Zhou et al., 2024] enhances the ZSAD performance of CLIP by proposing object-agnostic prompt learning. To tackle the issue of non-adaptation to common anomaly patterns, AdaCLIP [Cao et al., 2024] proposes hybrid dynamic learnable prompts which could flexibly adapt to the query images. However, existing ZSAD methods [Gu et al., 2024, Cao et al., 2023c, Oquab et al., 2023, Kirillov et al., 2023] cannot extract the generalizable anomaly patterns due to the object-level prompts. In contrast, our proposed HSP module could extract detailed features from the query image and learn semantic-rich information from type-level features, enhancing its sensitivity to anomalous features.
@@ 3.1 MulA: Multi-type Anomaly Dataset
Existing AD datasets mainly focus on providing object-level (e.g., “wood", “leather", “plastic") information [Bergmann et al., 2019a, Zou et al., 2022, Jezek et al., 2021, Head-CT-hemorrhage], while the type-level feature (e.g., “crack", “hole", “scratch") plays an essential role in extracting defect-region features from the query image. To this end, we propose Multi-type Anomaly dataset (MulA), which contains 11,226 images across 26 categories, surpassing existing AD datasets in scale and diversity. We collect the data from different scenarios, including industry, ordinary, and workshops, bringing multi-type generalizable anomaly patterns.
To build a dataset suitable for training, we manually annotated each sample with a high-quality mask. While some scenarios lack a sufficiently diverse set of defect types, others suffer from a shortage of normal (defect-free) samples. To address the former, we introduce Gaussian noise to enrich the defect distribution. For the latter, existing datasets often replicate a small number of normal samples, which leads to overfitting and hinders generalization in ZSAD. Therefore, we utilize image composition and geometric transformation to process different anomaly samples, obtaining normal samples with significant diversity. As shown in Figure 2(A) , MulA incorporates data across various scales, encompassing grayscale, RGB, and X-ray, exhibiting a wider generalization in industry. Besides the treatment of samples, we categorize industrial scenarios according to industry types. Furthermore, we compared classical datasets from both medical and industrial fields, as depicted in Table 1, thus validating the effectiveness of MulA. A detailed introduction to MulA can be found in Appendix Section 1.
Dataset Number Categories Modal
Visa 10,821 12 RGB
MVTec 5,354 15 RGB
MPDD 6,572 4 RGB
BTAD 2,830 3 RGB
DAGM 6,124 6 RGB
DTD 2,830 3 RGB
MulA 11,226 26 X-ray Grey RGB
Table 1: Comparison of MulA with large-scale industrial anomaly detection dataset.
Figure 2: (A) Statistics of the MulA Dataset and (B) examples of SEEK-M&V Knowledge Base.
@@ 3.2 SEEK-M&V: IAD Knowledge Base
Over the years, there is a lack of knowledge base for RAG frameworks in the field of anomaly detection [Zhao et al., 2024, Wang et al., 2025, Wu et al., 2025, Lee et al., 2025]. Moreover, current knowledge base tends to focus on textual information, while image features play an essential role in anomaly detection [Yu et al., 2024]. Building upon these insights, we build the first visual document knowledge base SEEK-M&V in IAD tasks, as illustrated in Figure 2(B).
To enhance the specialization of domain knowledge in the SEEK-M&V, we present detailed information about the anomaly with the official documentation for some of the AD datasets, such as anomaly type, anomaly analysis, and anomaly description. Meanwhile, we find production scenarios and application scenarios for various products to enrich the background information. Our SEEK-M&V contains both textual expertise and visual features. With generalizable forms, multimodal features and semantic-rich information, our SEEK-M&V has great potential in generalizing to other domains.
@@ 4 Method
In this section, we describe our proposed ADSeeker framework for anomaly detection and reasoning tasks, as shown in Figure 3. Recent studies reveal that pretrained MLLMs function as generalists, possessing a broad knowledge but underperforming in specialized domains, such as IAD. Therefore, our goal is to introduce a knowledge-injected framework designed to equip the generalist with the AD domain knowledge. This approach circumvents the need for pre-training while preserving the generalization capacity of the original model. We utilize the publicly available MLLM (Qwen2.5-VL [Bai et al., 2025] ) as the backbone of ADSeeker and the CLIP model (VIT-L/14@336px) as the encoder of the AD Expert module. The CLIP encoder aligns the visual information and the textual information to perform anomaly detection, and the LLM is responsible for textual instruction processing and complex reasoning.
Figure 3: The architecture of ADSeeker. It consists of two main knowledge-infused pathways: (1) The query image and knowledge documents are fed to the Q2K RAG to retrieve high-relevant domain knowledge. (2) The AD Expert integrates defect-region information and type-level features into semantic-rich visual tokens which will be passed into LLM.
@@ 4.1 Framework Overview
The framework of ADSeeker is illustrated in Figure 3. We store the type-level feature in learnable textual prompts $\left(T_{p},T_{n}\right)$, such as “A image with [cls] defect type”, which will be fed to CLIP Text Encoder. Similar to AdaCLIP, we leverage CLIP encoder to align the semantic-visual features. In the process of feature fusion, the visual embedding $E_{p}$ and textual embedding $E_{t}$ extracted by the CLIP Encoder will be fed to AD Expert module and then integrated into the anomaly prior feature $E_{c}$. Moreover, we introduce the HSP mechanism to preserve crucial visual features in CLIP. Specifically, learnable embeddings are leveraged to sparsify invalid information of query images layer by layer, while they will be combined with standard tokens at the beginning of each stage.
@@
The key features of MulA datasets undergo centroid-based clustering aligned with predefined cluster centers. Subsequently, key feature $K_{Q}$ is matched to the optimal cluster centroid to generate learnable textual prompts $T_{p},T_{n}$. Thanks to the high-quality data and type-level features of MulA dataset, our model could automatically generate accurate prompts and understand common defects among abnormal samples. Compared to commonly used textual prompts in ZSAD tasks, like “ A photo of a [obj]”, type-level prompts are generally more detailed and versatile in scope, as they encompass condensed content.
@@ 4.2 AD Expert:Hybrid Visual Fusion
As illustrated in Figure 3, we set a neural network to convert query image $I_{Q}$ into visual embeddings $Vis\in{R}^{x_{2}\times d}$, which will be integrated with the anomaly information to form anomaly prior embeddings $E_{c}={\{Loc,Vis\}}$. Given the semantic-rich visual embeddings, the defect characteristics will help the model generate fine-grained anomaly descriptions and enhance the image-level anomaly detection performance.
@@ 4.3 Q2K RAG:KDE-Sampled Multimodal RAG
Then the multimodal similarity $\mathcal{S}$ will be sorted to seek for the required domain knowledge $L_{a}ns$. As shown in Figure 4, in an ideal situation, the most relevant knowledge is most inspiring for the problem. We just need to retrieve the $\text{top-K}\downarrow$ expertise. Extensive experiments demonstrate that our seek accuracy rate has reached 83%. However, there is still some possibility that required domain knowledge $L_{a}ns$ is only at a relatively high rank, which greatly increases the difficulty of multimodal retrieval. Furthermore, SEEK-M&V is a knowledge base organized by categories, where different defect types of the same kind of object have a high degree of similarity. The simple top-k strategy has difficulty in retrieving product-related scenario information.

### v2

# ADSeeker: A Knowledge-Grounded Reasoning Framework for Industry Anomaly Detection and Reasoning

@@ Abstract
Automatic vision inspection holds significant importance in industry inspection. While multimodal large language models (MLLMs) exhibit strong language understanding capabilities and hold promise for this task, their performance remains significantly inferior to that of human experts. In this context, we identify two key challenges: (i) insufficient integration of anomaly detection (AD) knowledge during pre-training, and (ii) the lack of technically precise and context-aware language generation for anomaly reasoning. To address these issues, we propose ADSeeker, an anomaly task assistant designed to enhance inspection performance through knowledge-grounded reasoning. ADSeeker first leverages a curated visual document knowledge base, SEEK-MVTec&VisA (SEEK-M&V), which we construct to address the limitations of existing resources that rely solely on unstructured text. SEEK-M&V includes semantic-rich descriptions and image-document pairs, enabling more comprehensive anomaly understanding. To effectively retrieve and utilize this knowledge, we introduce the Query Image-Knowledge Retrieval-Augmented Generation (Q2K RAG) framework. To further enhance the performance in zero-shot anomaly detection (ZSAD), ADSeeker leverages the Hierarchical Sparse Prompt mechanism and type-level features to efficiently extract anomaly patterns. Furthermore, to tackle the challenge of limited industry anomaly detection (IAD) data, we introduce the largest-scale AD dataset, Multi-type Anomaly (MulA), encompassing 72 multi-scale defect types across 26 categories. Extensive experiments show that our plug-and-play framework, ADSeeker, achieves state-of-the-art zero-shot performance on several benchmark datasets.
@@ 1 Introduction
Figure 1: Visualization of the reasoning capability. Comparative results of ADSeeker against state-of-the-art models on the MMAD benchmark. Notably, the framework is specifically designed for IAD, its performance on defect analysis and localization tasks is better than any other IAD models.
@@
Figure 2: Industrial and medical image anomaly reasoning results of ADSeeker. The results include accurate anomaly location, anomaly description, defect classification, and other fine-grained information.
@@
Multimodal Large Language Models (MLLMs), such as GPT-4 and LLaVA [21], have shown significant promise across various intellectual human tasks, including writing, programming, and Industry Anomaly Detection (IAD). In IAD tasks, anomaly reasoning is distinct from anomaly detection. Anomaly reasoning necessitates a finer level of granularity and logical reasoning in the model’s responses. This distinction is vital in identifying the roles of a quality inspector versus a domain expert. Recent studies have explored the use of MLLMs for IAD tasks, including detection, description, analysis, etc., by training models such as AnomalyGPT [12] and Anomaly-OV [34], to incorporate domain-specific knowledge for improved anomaly reasoning. While some models show promise in Zero-Shot Anomaly Detection (ZSAD), they often falter in accurate defect analysis and detailed descriptions. Moreover, standard fine-tuning risks catastrophic forgetting of foundational alignments, compounded by limited datasets that cause overfitting to narrow answer templates and poor generalization, rather than robust context-aware anomaly reasoning. Tracing these performance bottlenecks to their core, we identify two key challenges that existing methods[37] have yet to fully address: i) insufficient integration of anomaly detection knowledge during pre-training, hindering precise localization, and ii) a lack of technically precise and context-aware language generation for anomaly reasoning. To tackle these issues, we propose ADSeeker, which addresses the limitations of domain knowledge through the Query Image—Knowledge Retrieval-Augmented-Generation (Q2K RAG).
@@
The Q2K RAG module achieves multimodal hybrid retrieval, seeking the knowledge document most relevant to the query image. Specifically, the query image and knowledge document are encoded into the Key Embedding and the Lock Embedding respectively. Subsequent multimodal retrieval is performed in the joint feature space between these representations. To compensate for the lack of knowledge base in IAD domain, we establish the first-ever knowledge base SEEK-M&V composed of visually rich industry anomaly documents. In SEEK-M&V, each knowledge document has specific reference pages containing the defect type and the defect analysis. Moreover, we utilize DeepSeek-R1 [10] to generate semantic-rich descriptions and application scenario introductions to expand the content of SEEK-M&V. As the saying goes, “a picture is worth a thousand words.” SEEK-M&V enriches knowledge representation by preserving graphic information for each anomaly type, offering a valuable reference for anomaly reasoning. Compared to existing training-based approaches[8], it significantly improves both efficiency and effectiveness.
@@
Existing approaches[22] have introduced learnable prompts to adapt models for ZSAD. However, object-level prompts (e.g., identifying objects such as “wood” or “leather”) are incapable of providing detailed anomaly patterns, and publicly available AD datasets struggle with semantically aligning these patterns. To address these gaps, we introduce Multi-type Anomaly (MulA), the largest-scale dataset for ZSAD, which includes 72 defect types across 8 scenarios. Additionally, ADSeeker utilizes a Hierarchical Sparse Prompt (HSP) mechanism to extract type-level features (e.g., distinguishing between “scratch” or “hole” anomalies) and reduce hallucinations caused by suspected anomaly regions.
@@
Extensive experiments validate the effectiveness and efficiency of ADSeeker. For example, anomaly reasoning experiments on MMAD [17] benchmark demonstrate that ADSeeker outperforms other general models. Additionally, experiments on 12 datasets from both industrial and medical sectors highlight ADSeeker’s superiority in ZSAD tasks. To validate the effectiveness of type-level features, we performed extensive tests to visualize the impact of type-level features across a range of products. We also conducted ablation studies and evaluated the performance-resource trade-offs within our ADSeeker.
@@
Figure 3: Details of MulA Dataset. (A) Statistics and components of the MulA Dataset, the newest dataset with the largest number and most types, and (B) examples of different types of type-level feature. Our dataset can also be organized by its defect type to enhance the Anomaly Detection performance of our ADSeeker through clustering.
@@ 2 Related Work
MLLM for industry anomaly detection (IAD). Recent advances have demonstrated the potential of MLLMs in IAD. For example, AnomalyGPT is directly fine-tuned on IAD datasets to inject domain knowledge into the base model. Moreover, to effectively extract detailed feature from abnormal samples, Myriad [23] developed an Expert Perception module, which is designed to harness the prior knowledge of vision experts. However, these models are constrained by the scope of their training knowledge and exhibit limited domain generalization. Existing IAD approaches[38] rarely become a trustworthy automatic vision inspection assistant [35, 9, 24, 25]. In contrast, building on the excellence of our ADSeeker framework, the anomaly reasoning performance of base models have been notably enhanced, as shown in Tab. 2.
Zero-Shot Anomaly Detection (AD). Recently, WinCLIP [15] makes an early attempt at introducing VLMs into ZSAD tasks, which leverage CLIP [28] to compute the similarities between patch embeddings and textual embeddings. On this basis, AnomalyCLIP [40] enhances the ZSAD performance of CLIP by proposing object-agnostic prompt learning. To tackle the issue of non-adaptation to common anomaly patterns, AdaCLIP [6] proposes hybrid dynamic learnable prompts which could flexibly adapt to the query images. However, existing ZSAD methods [11, 5, 27, 18] cannot extract the generalizable anomaly patterns due to the object-level prompts. In contrast, our proposed HSP module could extract detailed features from the query image and learn semantic-rich information from type-level features, enhancing its sensitivity to anomalous features at Fig. 3(B).
@@ 3.1 MulA: Multi-type Anomaly Dataset
Existing AD datasets mainly focus on providing object-level (e.g., “wood”, “leather”, “plastic”) information [4, 41, 16, 19], while the type-level feature (e.g., “crack”, “hole”, “scratch”) plays an essential role in extracting defect-region features from the query image. To this end, we propose Multi-type Anomaly dataset (MulA), which contains 11,226 images across 26 categories, surpassing existing AD datasets in scale and diversity. We collect the data from different scenarios, including industry, ordinary, and workshops, bringing multi-type generalizable anomaly patterns.
We constructed a high-quality training dataset through manual mask annotation. This dataset addresses two primary challenges: insufficient defect type diversity and a scarcity of normal (defect-free) samples. To enrich the defect distribution, we introduced Gaussian noise. Furthermore, to mitigate overfitting and poor ZSAD generalization caused by replicating limited normal samples, we generated a diverse set of normal samples by applying image composition and geometric transformations to existing anomaly data. As shown in Fig. 3(A) , MulA incorporates data across various scales, encompassing grayscale, RGB, and X-ray, exhibiting a wider generalization in industry. Besides the treatment of samples, we categorize industrial scenarios according to industry types. Furthermore, we compared classical datasets from both medical and industrial fields, as depicted in Fig. 3(A).
@@ 3.2 SEEK-M&V: IAD Knowledge Base
Over the years, there is a lack of knowledge base for RAG frameworks in the field of anomaly detection [39, 31, 32, 20]. Moreover, current knowledge base tends to focus on textual information, while image features play an essential role in anomaly detection [36]. Building upon these insights, we build the first visual document knowledge base SEEK-M&V in IAD tasks.
Figure 4: The architecture of ADSeeker. It consists of two main knowledge-infused pathways: (1) The query image and knowledge documents are fed to the Q2K RAG to retrieve high-relevant domain knowledge. (2) The AD Expert integrates defect-region information and type-level features into semantic-rich visual tokens which will be passed into VLM.
To enhance the specialization of domain knowledge in the SEEK-M&V, we present detailed information about the anomaly with the official documentation for AD datasets, such as anomaly type, anomaly analysis, and anomaly description. Meanwhile, we find production scenarios and application scenarios for various products to enrich the background information. Our SEEK-M&V contains both textual expertise and visual features. With generalizable forms, multimodal features and semantic-rich information, our SEEK-M&V has great potential in generalizing to other domains. These inherent strengths make it an effective solution to cross-disciplinary problems.
@@ 4 Method
In this section, we describe our proposed ADSeeker framework for anomaly detection and reasoning tasks, as shown in Fig. 4. Recent studies reveal that pretrained MLLMs function as generalists, possessing a broad knowledge but underperforming in specialized domains, such as IAD. Therefore, our goal is to introduce a knowledge-injected framework designed to equip the generalist with the AD domain knowledge. This approach circumvents the need for pre-training while preserving the generalization capacity of the original model. We utilize the publicly available MLLM (Qwen2.5-VL [3] ) as the backbone of ADSeeker and the CLIP model (VIT-L/14@336px) as the encoder of the AD Expert module. The CLIP encoder aligns the visual information and the textual information to perform anomaly detection, and the VLM is responsible for textual instruction processing and complex reasoning.
@@ 4.1 Framework Overview
The framework of ADSeeker is illustrated in Fig. 4(A). We store the type-level feature in learnable textual prompts $\left(T_{p},T_{n}\right)$, such as “A image with [cls] defect type”, which will be fed to CLIP Text Encoder. We leverage the CLIP encoder to align the semantic-visual features. In the process of feature fusion, the visual embedding $E_{p}$ and textual embedding $E_{t}$ extracted by the CLIP Encoder will be fed to AD Expert module and then integrated into the anomaly prior feature $E_{c}$. Moreover, we introduce the HSP mechanism to preserve crucial visual features in CLIP. Specifically, learnable embeddings are leveraged to sparsify invalid information from query images layer by layer, while they are combined with standard tokens at the beginning of each stage.
@@
The key features of MulA datasets undergo centroid-based clustering aligned with predefined cluster centers. Subsequently, key feature $K_{Q}$ is matched to the optimal cluster centroid to generate learnable textual prompts $T_{p},T_{n}$. With the type-level features of MulA dataset, our model could automatically generate accurate prompts and understand common defects among abnormal samples. Compared to commonly used textual prompts in ZSAD tasks, like “ Img of a [obj]”, type-level prompts are generally more detailed and versatile in scope, as they encompass condensed content.
@@ 4.2 AD Expert:Hybrid Visual Fusion
As illustrated in Fig. 4, we set a neural network to convert query image $I_{Q}$ into visual embeddings $Vis\in{R}^{x_{2}\times d}$, which will be integrated with the anomaly information to form anomaly prior embeddings $E_{c}={\{Loc,Vis\}}$. Given the semantic-rich visual embeddings, the defect characteristics will help the model generate fine-grained anomaly descriptions and enhance the image-level anomaly detection performance.
Figure 5: Illustration of the core architecture of Q2K RAG. Solving the problem of repetitive retrieval in industrial knowledge base through KDE-Sample strategy.
@@ 4.3 Q2K RAG:KDE-Sampled Multimodal RAG
Then the multimodal similarity $\mathcal{S}$ will be sorted to seek for the required domain knowledge $L_{a}ns$. As shown in Fig. 5, in an ideal situation, the most relevant knowledge is most inspiring for the problem. We just need to retrieve the $\text{top-K}\downarrow$ expertise. Extensive experiments demonstrate that our seek accuracy rate has reached 83%. However, there is still some possibility that required domain knowledge $L_{a}ns$ is only at a relatively high rank, which greatly increases the difficulty of multimodal retrieval. Furthermore, SEEK-M&V is a knowledge base organized by categories, where different defect types of the same kind of object have a high degree of similarity. The simple top-k strategy has difficulty in retrieving product-related scenario information.